Required information A friend tells you she has saved for seven years and has a present sum of $30,000, which she earned at the rate of 9% per year, compounded weekly. Determine the equivalent amount she had to start with seven years ago. The equivalent amount she had to start with seven years ago is?
Answer #1

i = 9% = 0.09 compounded weekly

i = 0.09 = 0.09 / 52 = 0.00173076 per week

Effective interest rate = ( 1+ 0.00173076)^52 - 1 = 0.094089165

t = 7 yrs

P = 30000 / (1+0.094089165)^7

P = 30000 / 1.876588378

P = 15986.46